Problem 10-10 (Algo) Multiple Products, Materials, and Processes [LO10-1, LO10-2] Mickley Corporation produces two products,Compute the materials price and quantity variances for each material. (Indicate the effect of each variance by selecting F
Answer #1

Solution:

Direct material variances - X442:

Actual price of material = $66,600 / 14800 = $4.50 per Kg

Material price variance = (SP -AP) *AQ Purchased = ($4.30 - $4.50) * 14800 = $2,960 U

Material Quantity variance = (SQ -AQ) *SP = [(2300*2 + 1600*3) - 9300] * $4.30 = $430 F

Direct material variances - Y661:

Actual price of material = $23,700 / 15800 = $1.50 per Liter

Material price variance = (SP -AP) *AQ Purchased = ($1.60 - $1.50) * 15800 = $1,580 F

Material Quantity variance = (SQ -AQ) *SP = [(2300*1.50 + 1600*2.50) - 13800] * $1.60 = $10,160 U